# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v2

@@ 5.2. Supporting Community Self-Governance in the Age of Generative AI
Comparing our findings to those from an earlier large-scale analysis of Reddit rules conducted by Fiesler et al. (Fiesler et al., 2018) provides insight into what is different about AI rules. We found that rules governing AI usage are more likely to be about images than the general rules analyzed in Fiesler et al. Together with the finding that images are the medium most referenced in AI rules, this finding suggests that the use of AI images may be the most common cause of community misalignment about AI norms. Our findings that Restrictive AI rules are more common than Prescriptive ones seems to suggest that, at least at this point in time, communities appear more interested in clarifying how AI cannot be used than how it can. Still, there is some evidence of flexibility in these stances, and it will be interesting to observe if community stances towards AI grow more or less flexible over time. It may be difficult to generalize about when communities require AI disclosure vs when they ban it outright, but digging into the motivations for these different approaches could be a topic of future HCI research.
@@
These rules—and norms—will continue to change over time. So far our data shows that AI rules have become more prevalent, but it is possible that these rules disappear if AI use becomes widespread to the point that it is unremarkable, or where bans have negative impacts on community engagement. Two of the top three most common rationales in the AI rules in our data were “Inauthentic” and “Low effort”, which seems to suggest that some communities perceive that AI use cheapens the social interactions that occur in online communities. These rationales for AI rules echo the findings from prior studies into the use of automated tools in interpersonal communication (Moradi and Levy, 2024; Rae, 2024). We suspect that even if communities cannot effectively ban AIGC, those that care about gate-keeping (Keegan and Gergle, 2010) may seek out other ways for users to demonstrate their authenticity, similar to those that have developed in some anonymous online communities (Bernstein et al., 2011). Platform designers can help by offering communication tools that encourage authenticity, such as those explored in recent work on effortful communication (Kelly et al., 2017; Zhang et al., 2022). Additionally, past HCI studies of online communities have shown that close-knit communities with strong social identity are able to navigate the challenging task of developing new norms around controversial technology use (Fiesler and Bruckman, 2019). In the age of generative AI, HCI research could develop novel ways to encourage strong social identity formation within online communities.
Our results complement Lloyd et al. (2023)’s qualitative investigation into moderators’ experience of enforcing AIGC rules by documenting the prevalence of these rules across Reddit. Their study identifies three main categories of moderator concern about how AIGC will impact their communities: concerns about content quality, social dynamics, and governance processes. Of these, content quality concerns are the most represented in our data. What a community considers to be “low quality” is extremely context-dependent, but this could be one area where UI features that clearly communicate norms and expectations to community members, such as rules that display at post creation time, could go a long way to positively shape user behavior (Matias, 2019). We suspect that rules about social dynamics are less prevalent in our data because of the phenomenon that we discuss in Section 5.1, where desired social dynamics may not need to be codified into rules because communities that care about them may be naturally better aligned. For example, in our data we do find rules about authenticity, but it tends to be more about the authenticity of content, not of interpersonal communication. Additionally, our findings provide further evidence of Lloyd et al.’s finding that moderators’ concerns about AIGC are based on their communities’ values. While rules usually do not explicitly state a community’s values, sometimes we can infer them, e.g. it is not a coincidence that Copyright and Piracy AI rules are more common in Art subreddits. Finally, we note that rules about images are very common in our data, while the majority of concerns raised in Lloyd et al. pertain primarily to text. This is a significant difference for the HCI research community to note, as helping communities respond to AI-generated text as opposed to AI-generated images is a different challenge with different potential solutions. For example, images may be more likely to be robustly watermarked (Dathathri et al., 2024) or to have provenance metadata (Feng et al., 2023), which platforms could clearly communicate to viewers.
@@
Our study is limited in that it is based on samples of subreddits that were publicly listed on Reddit’s web interface in November 2024 and July 2023. Our sample is thus missing subreddits that are public but unlisted, which makes it likely that certain types of subreddit are underrepresented, such as NSFW subreddits. Our sample also excluded subreddits with a primary language other than English. Future research could explore whether the trends that we observed still hold across the types of communities excluded from our sample. Our study also limits our analysis to text that is contained within a subreddit’s ”rules” metadata field. There are various other ways that online communities communicate rules and norms, such as through public discussions, wikis, direct messages, or moderation actions. Future work interested in studying communities’ attitudes towards AI use could consider these other sources of data, as well as qualitative approaches that would elicit direct input from community members and moderators.
@@
Ethics Statement. Our study was not required to undergo IRB review as we worked only with publicly available data. Still, we recognize the importance of conducting ethical social media research and have taken steps to follow the best practices of the HCI community (Fiesler et al., 2024). We did not collect any information about individual users or conversations. Additionally, we did not include in our study any subreddits who were not listed in Reddit’s public index of communities. We plan to make our datasets public after this publication completes the peer-review process, but will take additional steps to ensure no sensitive information is contained before doing so.
@@ 6. Conclusion
Communities across Reddit have responded to the arrival of generative AI by updating their public community rules with explicit positions on the use of this new technology. A specific community’s attitude towards the use of AIGC is nuanced and context-dependent, but our findings suggest a general trend of opposition from content generation communities that value quality and authenticity. It remains to be seen how these rules will be enforced, but HCI researchers should stay attuned to the changing needs and norms of these online communities as they continue to adapt to this new technology. Especially as social platforms begin incorporating generative AI features, it is important for designers to consider how to encourage the self-determination of communities with nuanced and varied stances on the use of AIGC.
@@ Acknowledgements.
Appendix A LLM Prompts
@@ A.2. Applying Rule Classification Labels
153Instructions: Does this rule ban or disallow AI or some sort og content made using AI without providing exceptions?

### v3

@@ 5.2. Supporting Community Self-Governance in the Age of Generative AI
Comparing our findings to those from an earlier large-scale analysis of Reddit rules conducted by Fiesler et al. (Fiesler et al., 2018) provides insight into what is different about AI rules. We found that rules governing AI usage are more likely to be about images than the general rules analyzed in Fiesler et al. Together with the finding that images are the medium most referenced in AI rules, this finding suggests that the use of AI images may be the most common cause of community misalignment about AI norms. Our findings that Restrictive AI rules are more common than Prescriptive ones seems to suggest that, at least at this point in time, communities appear more interested in clarifying how AI cannot be used than how it can. Still, there is some evidence of flexibility in these stances, and it will be interesting to observe if community stances towards AI grow more or less flexible over time.
@@
These rules—and norms about AI use—will continue to change over time. Past HCI studies have shown that a strong social identity can help close-knit online communities navigate the challenging task of developing new norms around controversial technology use (Fiesler and Bruckman, 2019), and we suspect the same will be true with generative AI. So far our data shows that AI rules have become more prevalent, but these rules may disappear if AI use becomes widespread to the point that it is unremarkable, or where bans have negative impacts on community engagement. Two of the top three most common rationales in the AI rules in our data were “Inauthentic” and “Low effort”, which seems to suggest that some communities perceive that the use of AI cheapens the social interactions that occur within them. These rationales for AI rules echo the findings from prior studies into the use of automated tools in interpersonal communication (Moradi and Levy, 2024; Rae, 2024). However, a closer look into our data reveals that these rules are more often about the authenticity of content, rather than interpersonal interactions. We will discuss these concerns about content authenticity more in the next paragraph. For communities who are concerned with authentic interactions, we suspect that even if communities cannot effectively ban AIGC, those that care about gate-keeping (Keegan and Gergle, 2010) will seek out other ways for users to demonstrate their authenticity, similar to those that have developed in some anonymous online communities (Bernstein et al., 2011). Platform designers can help by offering communication tools that encourage authenticity, maybe borrowing from the above, as well as from recent work on effortful communication (Kelly et al., 2017; Zhang et al., 2022).
Our results complement our earlier qualitative investigation into moderators’ experience of enforcing AIGC rules (Lloyd et al., 2024). That prior qualitative, interview-based study identified three main categories of moderator concern about how AIGC will impact their communities: concerns about content quality, social dynamics, and governance processes. Of these three concerns, the platform-wide quantitative analysis in this current work shows that concerns about content quality are the most prevalent in subreddit community rules. Our results in Section 4.2.2, where we found that AI rules are more common in Content Generation, Art, and Celebrity subreddits, suggest that rules about content quality might be most useful as ways to maintain standards in content-sharing communities. These communities tend to be less personal, with looser ties between members. We suspect that rules containing social dynamics concerns, highlighted in the earlier work (Lloyd et al., 2024), are less prevalent in our data because of the phenomenon that we discuss in Section 5.1, where desired social dynamics may not need to be codified into rules because the communities that care about them may be naturally better aligned.
Additionally, our quantitative results provide further evidence of the qualitative finding that moderators’ concerns about AIGC are based on their communities’ values (Lloyd et al., 2024). While rules usually do not explicitly state a community’s values, we can often infer them. For example, it is not a coincidence that Copyright and Piracy AI rules are more common in Art subreddits.
Finally, we note one interesting difference between the findings in our current study and those from our earlier qualitative work (Lloyd et al., 2024). While the majority of concerns raised in moderator interviews pertain primarily to text (Lloyd et al., 2024), rules about images are more common in our data. Helping communities respond to AI-generated text as opposed to AI-generated images is a different challenge with different potential solutions. For example, images may be more likely to be robustly watermarked (Dathathri et al., 2024) or to have provenance metadata (Feng et al., 2023), which platforms could clearly communicate to viewers.
@@
This current work is not without limitations. One limitation is that this study is based on a set of subreddits that were publicly listed on Reddit’s web interface in November 2024 (and July 2023). Our sample is thus missing subreddits that are public but unlisted, which makes it likely that certain types of subreddit are underrepresented, such as NSFW subreddits. Our sample also excluded subreddits with a primary language other than English. Future research could explore whether the trends that we observed still hold across the types of communities excluded from our sample. Our study also limits our analysis to text that is contained within a subreddit’s “rules” metadata field. There are various other ways that online communities communicate rules and norms, such as through public discussions, wikis, direct messages, or moderation actions. Future work interested in studying communities’ attitudes towards AI use could consider these other sources of data, as well as qualitative approaches that would elicit direct input from community members and moderators.
@@
Ethics Statement. Our study was not required to undergo IRB review as we worked only with publicly available data. Still, we recognize the importance of conducting ethical social media research and have taken steps to follow the best practices of the HCI community (Fiesler et al., 2024). We did not collect any information about individual users or conversations, only community-level metadata. Additionally, we did not include in our study any subreddits who were not listed in Reddit’s public index of communities. We make our datasets public with the publication of this work.
@@ 6. Conclusion
Communities across Reddit have responded to the arrival of generative AI by updating their public community rules with explicit positions on the use of this new technology. A specific community’s attitude towards the use of AIGC is nuanced and context-dependent, but our findings suggest a general trend of opposition, especially from content generation communities that value quality and authenticity. It remains to be seen how these rules will be enforced, but HCI researchers should stay attuned to the changing needs and norms of these online communities as they continue to adapt to this new technology. Especially as social platforms begin incorporating generative AI features, it is important for designers to consider how to encourage the self-determination of communities with nuanced and varied stances on the use of AIGC.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation CISE Graduate Fellowships under Grant No. 2313998, as well as Grant No. CHS 1901151/1901329. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Appendix A Appendix: LLM Prompts
@@ A.2. Applying Rule Classification Labels
153Instructions: Does this rule ban or disallow AI or some sort of content made using AI without providing exceptions?